# Harder-Narasimhan filtrations of decorated vector bundles

### v1

@@ 4. Harder-Narasimhan types
In this section, we learn how canonical reductions can be classified by obstruction classes and Harder-Narasimhan types, following [friedman-morgan_ontheconversetoatheoremofatiyahbott]. We then determine Harder-Narasimhan types for principal bundles with structure groups $\mathbf{GL}(r,\mathbb{C})$, $\mathbf{SO}(r,\mathbb{C})$, and $\mathbf{Sp}(2n,\mathbb{C})$. To do this, we first review real forms $K$ of complex reductive groups $G$, which are maximal compact subgroups of $G$. Then, we review the algebraic characterization of the fundamental groups of $T$, $G$, $[G,G]$ and $G_{\mathrm{ab}}=G/[G,G]$ as $\mathbb{Z}$-modules. Using Čech cohomology, we are then able to define obstruction classes and topological types of principal bundles.
@@ 4.1. Fundamental groups of reductive groups
We follow [fulton-harris_representationtheory, §26.1] to investigate real forms of $G$ and $\mathrm{Lie}(G)$.
@@
We first assume that $G$ is semisimple. By [fulton-harris_representationtheory, Proposition 26.4] and its following discussion, we know that there exists a unique real form $\underline{k}$ of $\mathrm{Lie}(G)$, up to inner automorphism, that induces a compact real Lie subgroup $K$ of $G$ (through the exponential map). To show that $K$ is also a compact real form, we note that $\underline{k}$ yields a decomposition $\mathrm{Lie}(G)\cong\underline{k}\oplus\underline{p}$ at the level of real Lie algebras, induced by a Cartan involution $\theta$, due to [knapp_liegroups, VI.2. Corollary 6.22]. From this, we can use [knapp_liegroups, Theorem 6.31 (g)] to follow that $K$ is a maximal compact subgroup of $G$, i.e., a compact real form. In the general case where $G$ reductive, we know that $[G,G]$ is semisimple, for which the above argument yields a compact real form $K^{\prime}$ of $[G,G]$, whose Lie algebra $\mathrm{Lie}(K^{\prime})$ is a compact real form of $\mathrm{Lie}([G,G])$. Since $G=[G,G]\mathbf{R}(G)$ due to [borel_linearalgebraicgroups, IV. 14.2 Proposition], and $\mathbf{R}(G)\cong(\mathbb{C}^{\times})^{q}$ is a torus due to [borel_linearalgebraicgroups, IV. 11.21 Proposition], we choose a compact real Lie subgroup $R\cong\mathbf{U}(1)^{q}$ of $\mathbf{R}(G)$, and define $K=K^{\prime}R$. It follows that $K$ is a compact real form of $G$, where $\mathrm{Lie}(K)$ is a compact real form of $\mathrm{Lie}(G)$. ∎
@@
Let us fix a Cartan subgroup $T\cong(\mathbb{C}^{\times})^{r}$ of $G$, and a real form $K$ of $G$. For $H=K\cap T$, we have that $H\cong\mathbf{U}(1)^{r}$ is a maximal compact torus of $K$. We want algebraic characterizations of the fundamental groups of $T$, $G$, $[G,G]$, and $G_{\mathrm{ab}}$, as $\mathbb{Z}$-modules. Some of these fundamental groups embed naturally as lattices of the real vector spaces, as defined in [friedman-morgan_ontheconversetoatheoremofatiyahbott, 1]:
@@
We now recall some constructions from abstract root systems, applied to the root system $\Phi(G,T)$ of $G$ on $(\mathbf{E},\langle\_,\_\rangle)$, the Euclidean space isomorphic to $\mathbb{R}^{r}$ where $\langle\_,\_\rangle$ is a nondegenerate symmetric form induced from the Killing-form in [humphreys_introductiontoliealgebras, II.8.5].
@@
By [*][III 9.2]humphreys_introductiontoliealgebras, it is known that $\Phi^{\vee}(G,T)$ is a root system of $(\mathbf{E},\langle\_,\_\rangle)$. It is also clear that dual-roots and dual-coroots form root systems.
@@
Note that $\mathbf{E}$ is a subspace of $\mathrm{Lie}(T)$. Furthermore, the dual-coroot lattice $\Lambda:=\mathrm{span}_{\mathbb{Z}}(\Phi_{H}^{\vee}(G,T))$ is contained within the kernel lattice $\Gamma$ from Remark 4.5, as proven in [*][Lemma 12.8]hall_liegroupsliealgebrasandreps.
We can now search for expressions of $\pi_{1}(G)$, $\pi_{1}([G,G])$, and $\pi_{1}(G_{\mathrm{ab}})$.
@@
For the proof of (i), see [*][Corollary 13.18]hall_liegroupsliealgebrasandreps, noting that $T\hookrightarrow G$ is the complexified version of $H\hookrightarrow K$. In particular, the surjectivity of $\pi_{1}(T)\rightarrow\pi_{1}(G)$ is proven in [*][Proposition 13.37]hall_liegroupsliealgebrasandreps.
We prove (ii). Since $G_{\mathrm{ab}}\cong(\mathbb{C}^{\times})^{q}$, we have that $\pi_{1}(G_{\mathrm{ab}})\cong\mathbb{Z}^{q}$ is a lattice. As $[G,G]$ has a semisimple real form $[K,K]$, Lie algebra cohomology implies that $\pi_{1}([G,G])$ is finite, as seen in [*][Theorem 16.1]chevalley-eilenberg_theoreyofliegroups.
@@
Since $[G,G]$ is connected, we have $\pi_{0}([G,G])\cong 1$. Using Morse theory results from [bott_morsetheorytopologyofliegroups], it can be proven that $\pi_{2}(G_{\mathrm{ab}})\cong 1$, as $G_{\mathrm{ab}}$ is a Lie group. Therefore, the sequence in (4.1) is exact. The isomorphisms $\pi_{1}(G_{\mathrm{ab}})\cong\Gamma/\widehat{\Lambda}$, and $\pi_{1}([G,G])\cong\widehat{\Lambda}/\Lambda$, follow directly from (4.1) and (i), (ii). ∎
@@
    $\mathrm{Hom}(\mathbf{z}(\mathrm{Lie}(K)),\mathrm{Lie}(\mathbf{U}(1)))=\mathrm{Hom}_{\mathbb{R}}(\mathbf{i}\mathbf{z}(\mathrm{Lie}(G))_{\mathbb{R}},\mathbf{i}\mathbb{R})\cong\mathrm{Hom}_{\mathbb{R}}(\mathbf{z}(\mathrm{Lie}(G))_{\mathbb{R}},\mathbb{R})=\mathbf{z}(\mathrm{Lie}(G))^{*}_{\mathbb{R}}.$
@@ 4.2. Obstruction classes and Harder-Narasimhan types
By construction, the sets $\check{H}^{1}(X,\mathbf{O}_{X}(\_))$ correspond to sets of isomorphism classes of principal bundles. These isomorphism classes can be represented by cocycles of principal bundles, which correspond to $1$-cocycles in the sense of Čech cohomology. As $X$ is a compact Riemann surface, the singular cohomologies $H^{*}(X,\pi_{1}(G))$ and Čech cohomologies $\check{H}^{*}(X,\mathbf{O}_{X}(\pi_{1}(G)))$ are naturally isomorphic, as seen in [ramanan_globalcalculus, Chapter 4. 4.14. Theorem, 5.3. Theorem]. Applying Poincaré duality then gives us a canonical isomorphism $\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))\cong\pi_{1}(G)$. We now investigate the connecting map $o_{2}:\check{H}^{1}(X,\mathbf{O}_{X}(G))\rightarrow\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))$. For a principal-$G$-bundle $\xi$, its isomorphism class $\xi]\in\check{H}^{1}(X,\mathbf{O}_{X}(G))$ is represented by the cocycles $(\sigma_{i,j})_{i,j\in J}$. For all $i,j\in J$, we use the universal lifting property of covering spaces to lift $\sigma_{i,j}$ to $\widetilde{\sigma_{i,j}}$, such that the following diagram commutes:
@@
This defines a $2$-cocycle $(\sigma_{i,j,k})_{i,j,k\in J}$ representing $o_{2}([\xi])\in\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))$. If $o_{2}([\xi])=1\in\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))$, then $\xi]$ is equivalently the image of an element $)[\widetilde{\xi}]\in\check{H}^{1}(X,\mathbf{O}_{X}(\widetilde{G}$)), i.e., there exists a principal-$\widetilde{G}-$bundle $\widetilde{\xi}$ such that $\widetilde{\xi}(G)\cong\xi$.
@@
[…]
@@
We are now able to state the existence and uniqueness of canonical reductions, as proven in Subsections 3.2 and 3.3, using types, as shown in [friedman-morgan_ontheconversetoatheoremofatiyahbott, Lemma 2.2.1].
@@
For a principal-$G$-bundle $\xi$, and a reduction $s^{*}\xi$ of $\xi$ as in Theorem 4.19, we call $\mu^{\xi}_{\mathrm{HN}}=\mu^{(s^{*}\xi)(L_{I})}\in\mathrm{Lie}(T)_{\mathbb{R}}$ the Harder-Narasimhan type of $\xi$.
@@
We fix a Cartan subgroup and Borel subgroup of $G$ such that $T\subseteq B\subseteq G$. The smooth moduli stack $Bun(G)$ of principal-$G$-bundles on a Riemann surface $X$, where $G$ is a complex reductive group, has a stratification:
@@
One can define a partial order $\geqslant$ on the pairs $(\mu,P_{I})$ of Harder-Narasimhan types $\mu$, together with their corresponding standard parabolics $P_{I}$ in the canonical reductions, as done in [friedman-morgan_ontheconversetoatheoremofatiyahbott, Definition 2.4.1]:
@@
where $\overline{Bun_{\mu}(G)}$ is the closure of $Bun_{\mu}(G)$ in $Bun(G)$. This stratification can be seen as a colimit of GIT-instability (Hesselink) stratifications, as described more generally for the stack of coherent sheaves in [hoskins_stratificationsofmoduliofsheaves, 4.4, 4.5, 4.6].

### v2

@@ 5. Harder-Narasimhan types
In this section, we learn how canonical reductions can be classified by obstruction classes and Harder-Narasimhan types, following [FM02] and [HL09]. We then determine Harder-Narasimhan types for principal bundles with structure groups $\mathbf{GL}(r,\mathbb{C})$, $\mathbf{SO}(r,\mathbb{C})$, and $\mathbf{Sp}(2n,\mathbb{C})$. To do this, we first review real forms $K$ of complex reductive groups $G$, which are maximal compact subgroups of $G$. Then, we review the algebraic characterization of the fundamental groups of $T$, $G$, $[G,G]$ and $G_{\mathrm{ab}}=G/[G,G]$ as $\mathbb{Z}$-modules. Using Čech cohomology, we are then able to define obstruction classes and topological types of principal bundles.
@@ 5.1. Fundamental groups of reductive groups
We follow [FH04, §26.1] to investigate real forms of $G$ and $\mathrm{Lie}(G)$.
@@
We first assume that $G$ is semisimple. By [FH04, Proposition 26.4] and its following discussion, we know that there exists a unique real form $\underline{k}$ of $\mathrm{Lie}(G)$, up to inner automorphism, that induces a compact real Lie subgroup $K$ of $G$ (through the exponential map). To show that $K$ is also a compact real form, we note that $\underline{k}$ yields a decomposition $\mathrm{Lie}(G)\cong\underline{k}\oplus\underline{p}$ at the level of real Lie algebras, induced by a Cartan involution $\theta$, due to [Kna88, VI.2. Corollary 6.22]. From this, we can use [Kna88, Theorem 6.31 (g)] to follow that $K$ is a maximal compact subgroup of $G$, i.e., a compact real form. In the general case where $G$ reductive, we know that $[G,G]$ is semisimple, for which the above argument yields a compact real form $K^{\prime}$ of $[G,G]$, whose Lie algebra $\mathrm{Lie}(K^{\prime})$ is a compact real form of $\mathrm{Lie}([G,G])$. Since $G=[G,G]\mathbf{R}(G)$ due to [Bor91, IV. 14.2 Proposition], and $\mathbf{R}(G)\cong(\mathbb{C}^{\times})^{q}$ is a torus due to [Bor91, IV. 11.21 Proposition], we choose a compact real Lie subgroup $R\cong\mathbf{U}(1)^{q}$ of $\mathbf{R}(G)$, and define $K=K^{\prime}R$. It follows that $K$ is a compact real form of $G$, where $\mathrm{Lie}(K)$ is a compact real form of $\mathrm{Lie}(G)$. ∎
@@
Let us fix a Cartan subgroup $T\cong(\mathbb{C}^{\times})^{r}$ of $G$, and a real form $K$ of $G$. For $H=K\cap T$, we have that $H\cong\mathbf{U}(1)^{r}$ is a maximal compact torus of $K$. We want algebraic characterizations of the fundamental groups of $T$, $G$, $[G,G]$, and $G_{\mathrm{ab}}$, as $\mathbb{Z}$-modules. Some of these fundamental groups embed naturally as lattices of the real vector spaces, as defined in [FM02, 1]:
@@
We now recall some constructions from abstract root systems, applied to the root system $\Phi(G,T)$ of $G$ on $(\mathbf{E},\langle\_,\_\rangle)$, the Euclidean space isomorphic to $\mathbb{R}^{r}$ where $\langle\_,\_\rangle$ is a nondegenerate symmetric form induced from the Killing-form in [Hum72, II.8.5].
@@
By [Hum72, III 9.2], it is known that $\Phi^{\vee}(G,T)$ is a root system of $(\mathbf{E},\langle\_,\_\rangle)$. It is also clear that dual-roots and dual-coroots form root systems.
@@
Note that $\mathbf{E}$ is a subspace of $\mathrm{Lie}(T)$. Furthermore, the dual-coroot lattice $\Lambda:=\mathrm{span}_{\mathbb{Z}}(\Phi_{H}^{\vee}(G,T))$ is contained within the kernel lattice $\Gamma$ from Remark 5.5, as proven in [Hal15, Lemma 12.8].
We can now search for expressions of $\pi_{1}(G)$, $\pi_{1}([G,G])$, and $\pi_{1}(G_{\mathrm{ab}})$, following [FM02, $\S$2] and [HL09, $\S$3.3].
@@
For the proof of (i), see [Hal15, Corollary 13.18], noting that $T\hookrightarrow G$ is the complexified version of $H\hookrightarrow K$. In particular, the surjectivity of $\pi_{1}(T)\rightarrow\pi_{1}(G)$ is proven in [Hal15, Proposition 13.37].
We prove (ii). Since $G_{\mathrm{ab}}\cong(\mathbb{C}^{\times})^{q}$, we have that $\pi_{1}(G_{\mathrm{ab}})\cong\mathbb{Z}^{q}$ is a lattice. As $[G,G]$ has a semisimple real form $[K,K]$, Lie algebra cohomology implies that $\pi_{1}([G,G])$ is finite, as seen in [CE48, Theorem 16.1].
@@
Since $[G,G]$ is connected, we have $\pi_{0}([G,G])\cong 1$. Using Morse theory results from [Bot56], it can be proven that $\pi_{2}(G_{\mathrm{ab}})\cong 1$, as $G_{\mathrm{ab}}$ is a Lie group. Therefore, the sequence in (5.1) is exact. The isomorphisms $\pi_{1}(G_{\mathrm{ab}})\cong\Gamma/\widehat{\Lambda}$, and $\pi_{1}([G,G])\cong\widehat{\Lambda}/\Lambda$, follow directly from (5.1) and (i), (ii). ∎
@@
    $\mathrm{Hom}(\mathbf{z}(\mathrm{Lie}(K)),\mathrm{Lie}(\mathbf{U}(1)))$ $=\mathrm{Hom}_{\mathbb{R}}(\mathbf{i}\mathbf{z}(\mathrm{Lie}(G))_{\mathbb{R}},\mathbf{i}\mathbb{R})$
    $\cong\mathrm{Hom}_{\mathbb{R}}(\mathbf{z}(\mathrm{Lie}(G))_{\mathbb{R}},\mathbb{R})=\mathbf{z}(\mathrm{Lie}(G))^{*}_{\mathbb{R}}.$
@@ 5.2. Obstruction classes and Harder-Narasimhan types
By construction, the sets $\check{H}^{1}(X,\mathbf{O}_{X}(\_))$ correspond to sets of isomorphism classes of principal bundles. These isomorphism classes can be represented by cocycles of principal bundles, which correspond to $1$-cocycles in the sense of Čech cohomology. As $X$ is a compact Riemann surface, the singular cohomologies $H^{*}(X,\pi_{1}(G))$ and Čech cohomologies $\check{H}^{*}(X,\mathbf{O}_{X}(\pi_{1}(G)))$ are naturally isomorphic, as seen in [Ram05, Chapter 4. 4.14. Theorem, Chapter 4. 5.3. Theorem]. Applying Poincaré duality then gives us a canonical isomorphism $\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))\cong\pi_{1}(G)$. We now investigate the connecting map $o_{2}:\check{H}^{1}(X,\mathbf{O}_{X}(G))\rightarrow\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))$. For a principal $G$-bundle $\xi$, its isomorphism class $\xi]\in\check{H}^{1}(X,\mathbf{O}_{X}(G))$ is represented by the cocycles $(\sigma_{i,j})_{i,j\in J}$. For all $i,j\in J$, we use the universal lifting property of covering spaces to lift $\sigma_{i,j}$ to $\widetilde{\sigma_{i,j}}$, such that the following diagram commutes:
@@
This defines a $2$-cocycle $(\sigma_{i,j,k})_{i,j,k\in J}$ representing $o_{2}([\xi])\in\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))$. If $o_{2}([\xi])=1\in\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))$, then $\xi]$ is equivalently the image of an element $)[\widetilde{\xi}]\in\check{H}^{1}(X,\mathbf{O}_{X}(\widetilde{G}$)), i.e., there exists a principal $\widetilde{G}-$bundle $\widetilde{\xi}$ such that $\widetilde{\xi}(G)\cong\xi$.
@@
Let $\xi$ be a principal $G$-bundle on $X$. Via the isomorphism $\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))\cong\pi_{1}(G)$, the class $o_{2}([\xi])\in\check{H}^{2}(X,\mathbf{O}_{X}(\pi_{1}(G)))$ induces an element $o_{2}(\xi)\in\pi_{1}(G)$, called the (second) obstruction class of $\xi$.
@@
We are now able to state the existence and uniqueness of canonical reductions, as proven in Subsections 4.2 and 4.3, using types, as shown in [FM02, Lemma 2.2.1].
@@
For a principal $G$-bundle $\xi$, and a reduction $s^{*}\xi$ of $\xi$ as in Theorem 5.19, we call $\mu^{\xi}_{\mathrm{HN}}=\mu^{(s^{*}\xi)(L_{I})}\in\mathrm{Lie}(T)_{\mathbb{R}}$ the Harder-Narasimhan type of $\xi$.
@@
We fix a Cartan subgroup and Borel subgroup of $G$ such that $T\subseteq B\subseteq G$. The smooth moduli stack $Bun(G)$ of principal $G$-bundles on a Riemann surface $X$, where $G$ is a complex reductive group, has a stratification:
@@
One can define a partial order $\geqslant$ on the pairs $(\mu,P_{I})$ of Harder-Narasimhan types $\mu$, together with their corresponding standard parabolics $P_{I}$ in the canonical reductions, as done in [FM02, Definition 2.4.1]:
@@
where $\overline{Bun_{\mu}(G)}$ is the closure of $Bun_{\mu}(G)$ in $Bun(G)$. This stratification can be seen as a colimit of GIT-instability (Hesselink) stratifications, as described more generally for the stack of coherent sheaves in [Hos18, 4.4, 4.5, 4.6].